OpenStudy (anonymous):

Consider the function f(x) = \frac{64 x^3 - 80 x^2+ 24 x}{4 x^3+ 13 x^2 - 32 x+ 15 } The root(s) of f(x) are: Enter your answer as a comma-separated list. The graph of f(x) has hole(s) when x has the following value(s): Enter your answer as a comma-separated list.

OpenStudy (ranga):

For the roots, the numerator has to be zero: 64x^3 - 80x^2+ 24x = 0 Solve for x. You can find the GCF and factor it out of the numerator. Then you will have a quadratic which you know how to factor or solve. Being a cubic equation you should get 3 roots. f(x) is undefined when the denominator is zero. Set 4 x^3+ 13 x^2 - 32 x+ 15 = 0 and solve for x to find the holes.

OpenStudy (anonymous):

I got three roots but they are wrong.

OpenStudy (ranga):

64x^3 - 80x^2+ 24x = 0 8x(8x^2 - 10x + 3) = 0 8x = 0 or (8x^2 - 10x + 3) = 0 x = 0 is one root 8x^2 - 10x + 3 = 0 8x^2 - 6x - 4x + 3 = 0 2x(4x - 3) - 1(4x - 3) = 0 (2x - 1)(4x - 3) = 0 x = 1/2 or 3/4 The roots are: x = 0, 1/2, 3/4

OpenStudy (anonymous):

That is what I got for the roots but it still says those are wrong :/ Thanks for the help.

OpenStudy (ranga):

Oh, I know why it says it is wrong. The mystery will be solved when you find the holes by setting the denominator to 0. For those x values f(x) will be undefined and one of the holes happens to be 3/4 which you have to drop it from the roots. So the roots will be just: 0 and 1/2.

OpenStudy (anonymous):

It worked! Thank you!

OpenStudy (ranga):

You are welcome. Don't forget to do the second part of the problem where you have to find the x values for which the function is not defined (f(x) has "holes"). You need to solve: 4 x^3+ 13 x^2 - 32 x+ 15 = 0